a manufacturer determines that the profit in dollars for manufacturing n units is P= 3n^2-30n-170. (Assume that n is a postive integer) How many units are produced when the profit is $430?
Answer by ewatrrr(24785)

 
Hi,
P= 3n^2-30n-170
P = $430
430 = 3n^2 - 30n - 170
OR
3n^2 -30n - 600 = 0
n^2 - 10n - 200 = 0
factoring
(n - 20)(n + 10) = 0 ||ony positive solution is appropriate
n = 20, number of units produced when the profit is $430